The molecule below contains carbon atoms (shown in
Answer #1
Concepts and reason

The 3D structure of a molecule is given and the different functional groups are needed to be given. Label the structure as per the notations given and convert the 3D structure to full displayed skeletal structure and find the groups from the structure obtained.

Fundamentals

Skeletal structure: The skeletal formula represents the bonding of different atoms with each other and shows the basic skeleton of the molecule.

The several groups that a molecule can possess are listed below:

• Alkyl halide: When a halogen is directly attached to carbon atom, then it is known as alkyl halide. They are represented as R-X, where R is an alkyl group and X represents halogen.

• Alcohol: The alcohol group is represented as -OH.

• Alkyne: when a triple bond is present between carbon atoms, then it is known as an alkyne.

• Acid anhydride: In an acid anhydride, two acyl groups are bonded with a common oxygen atom. It is represented as follows:

R
Anhydride group

Where R and R’ are different alkyl groups.

• Thiol: It is represented as -SH.

• Amine: It is represented as NH2 - {\rm{N}}{{\rm{H}}_2} .

• Carboxylic acid: It is represented as -COOH. It is shown as follows:

R
НО
H
Carboxylic acid

Where R is an alkyl group.

• Aldehyde and ketone: Both the groups have a carbonyl group (C=O)\left( {{\rm{C = O}}} \right) . When the carbonyl group is attached to one alkyl group, then it is known as aldehyde and when both the groups attached are alkyl groups, then it is known as ketone.

R
R
R
Aldehyde
Ketone

Where R, R’, R” are the different alkyl groups.

• Amide: It is represented as CONH2 - {\rm{CON}}{{\rm{H}}_2} . It is shown below:

NH2
R
Amide

• Ether: When an oxygen atom is attached to two alkyl groups, then it is known as ether group. It is represented as R-O-R’, where R and R’ are different alkyl groups.

• Ester: it is represented as -COOR. It is shown below:

OR
R
Ester

• Alkene: when a double bond is present between carbon atoms, then it is known as an alkene.

• Acid chloride: It is represented as -COCl. It is shown below:

Cl
R
Acid chloride

• Arene/aromatic: These groups have (4n+2)π\left( {4n + 2} \right)\pi electrons and consist of a ring where all the constituent atoms are present in a plane. An example is given below:

6 pi electrons

Label the different colored spheres according to the atoms mentioned. The black spheres represent carbon, white represent hydrogen and red represent oxygen.

Н
С
С
-С.
н
-С.
н
н
Н—с.
Н
Н
Н

The various groups present in the molecule are ketone, ether and alkyne. They are represented as follows:

Н
Ketone
Alkyne
Н
н—с
н
Н
Ether

Ans:

The groups present in the given molecule are ketone, ether and alkyne.